# arrows.meta, chains, positioning

### v1

@@ IV-C GraphRL Agent
In this study, the game learning agent Deep Q-Networks (DQN) algorithm is used. The algorithm was introduced by Google’s DeepMind for playing Atari game to play games by just observing the screen without any training or prior knowledge about those games. In this algorithm, the Q-Learning functions’ approximation is computed using the proposed T-GCN model and the learning agent gets rewarded based on the graph network prediction of the right action for the current state.
@@ IV-C1 Q-Function Approximation
T-GCN model used in this study to approximate the Q-Function for each action in the action space as shown in Fig. 1. The model is configured with parameters such as the relu activation function, mean square error as loss function, and Adam optimiser. The model gets trained with the state and its corresponding action. The learning agent performs an action $a_{t}\epsilon A$ for a transition from state $s_{t}$ to $s_{t}^{{}^{\prime}}$ and achieves a reward $R$ for the action. In this transition process, the maximum of the Q-function in Equation 7 is calculated, and the discount of the calculated value uses a discount factor $\gamma$ to suppress future rewards and focus on immediate rewards. The discounted future reward is added to the current reward to get the target value. The difference between the current prediction from the neural networks and the calculated target value provides a loss function. The loss function is a deviation of the predicted value from the target value and it can be estimated from Equation 8. The square of the loss function allows for the punishment of the agent for a large loss value.
@@ IV-C3 GraphRL Agent Algorithm
4:if $np.random.rand<\epsilon$ then $\triangleright$ Exploration
@@
13:         $target\leftarrow R+\gamma(max(model.predict(s_{t+1})))$
@@
19:if $\epsilon$geq $\epsilon_{min}$ then
20:     $\epsilon$= $\epsilon_{decay}$
@@
The Learning Agent algorithm 2 encapsulates methods for function approximation, memorization, replay, exploration, and exploitation. In lines 1-2, hyperparameters required for the agent and T-GCN model for Q-function approximation are initialized. Memorization and replay, explained in line 3, involve storing T-GCN experience, including state, action, reward, and next_state, for later retraining of the model using the replay method. The algorithm (lines 4-8) predicts an action using either exploration or exploitation methods. Finally, lines 9-21 retrieve a batch of previous experiences from memory, process and retrain the neural network model based on the earlier defined hyperparameters.
@@ IV-C4 Implementation Algorithm
2:$\mathcal{C}={1,2,\dots,C}$: set of subjects
3:$\mathcal{V}={1,2,\dots,V}$: set of vital signs
4:$\mathcal{M}={1,2,\dots,M}$: number of episodes
@@
The Algorithm  3 corresponds to the previous two algorithms,  1 and  2, and implements the proposed GraphRL monitoring framework. It takes as input a set of subjects with their vital signs and the number of training episodes for the agents. The output of the algorithm is the cumulative reward score achieved by the learning agent in each episode. The first two lines of the algorithm create instances of the Predictive GraphRL environment and GraphRL agent. Lines 3-17 consist of nested for loops with conditional statements to determine if an episode is complete. The outer loop iterates through each episode, resetting the environment to its initial state and the score to zero. The inner loop iterates through timesteps, which represent the current state’s time, and calls the methods defined in Algorithms  1 and  2 to predict the current state’s action, reward the agent for its predicted action, retrieve the next state, and store previous experiences. Finally, the replay method is called to update the neural network model with the stored experiences.
@@ IV-D Bayesian Optimisation
Bayesian optimisation is a global optimisation method that uses a probabilistic model to guide the search for the optimal solution. The model is updated as new data points are sampled and evaluated, allowing the algorithm to improve its predictions over time by fine-tuning the $L_{reg}$, $\lambda$ and minimising the loss function defined in Equation 6. The basic idea behind Bayesian optimisation is to model the objective function, f(x), as a Gaussian process (GP). The GP model is used to predict the objective function value at any point x, given the observations of the function at other points. The prediction is given by the posterior distribution of the GP, which is a Gaussian distribution with mean and variance given by Equation 9.
@@
Where X is the matrix of previously sampled points, y is the vector of corresponding function values, K is the Gram matrix of the covariance function evaluated at X, and $\sigma^{2}$ is the noise level in the function evaluations. The next point to sample is chosen based on an acquisition function, which balances the trade-off between exploration and exploitation. Common acquisition functions include the probability of improvement and the expected improvement. Given a set of observed points (X,y) and a Gaussian process prior, Bayesian optimisation seeks the point x* that minimises the loss function value, given by Equation 10. The optimisation process continues iteratively, sampling new points and updating the GP model until a stopping criterion is met.
@@ V-A Datasets
In this study, the proposed framework is evaluated with three publicly available datasets related to healthcare, traffic and weather.
  • Wearable Stress and Affect Detection (WESAD) dataset [32]: This is a publicly available dataset that contains physiological and self-report data collected from 17 participants during various stress and affective states.The dataset contains data collected from various wearable sensors, including a chest-mounted electrocardiogram (ECG) sensor, a wrist-mounted photoplethysmogram (PPG) sensor, and an accelerometer.
  • Los Angeles (LA) traffic dataset [33]: This is a publicly available dataset that contains data on traffic flow in the Los Angeles metropolitan area. The dataset was collected by the Los Angeles Department of Transportation (LADOT) and contains information on traffic counts, traffic speeds, and travel times for a variety of purposes, including traffic prediction, traffic flow analysis, and transportation planning.
  • Large-Scale Traffic and Weather Events (LSTW) dataset [34]: It contains information on traffic conditions and extreme weather events in the United States. In terms of weather events, the dataset has several types including rain, snow, storm, cold weather event, etc.
@@ V-B Baseline Models
  • In Li et al.’s work [35], the authors suggested the ELMA method for multi-agent activity forecasting through graph neural networks that incorporate spatiotemporal data using simultaneous energy-based learning.
  • Ma et al. [36] proposed a technique for predicting multi-agent driving behaviour in different scenarios based on self-supervised domain knowledge.
  • In Jiang et al.’s research [37], they conducted a study on internet traffic prediction with distributed multi-agent learning and utilized both LSTM and gated recurrent unit (GRU). Their GRU-based distributed multi-agent learning scheme outperformed LSTM.
@@ V-C Evaluation Metrics
Mean Absolute Error (MAE) is a commonly used regression metric that determines the average magnitude of errors between the predicted and actual values in a given dataset. It is calculated by taking the average of the absolute differences between the predicted and actual values, which results in a single value. Another popular regression metric is the Root Mean Squared Error (RMSE), which measures the average magnitude of the differences between the predicted and actual values. RMSE is calculated as the square root of the mean of the squared differences between the predicted and actual values. Mean Absolute Percentage Error (MAPE) is another regression metric that calculates the average absolute percentage error between the predicted and actual values. It is calculated as the average of the absolute differences between the predicted and actual values, expressed as a percentage of the actual values. Cumulative Rewards, on the other hand, is a performance metric used in reinforcement learning that quantifies the total rewards obtained by an agent within a specific time frame or number of actions. It is computed by summing up all the rewards that the agent received during the given period or a number of actions. The experiments conducted in this study employed Python version 3.7.6 and utilized several packages, including TensorFlow, Keras, Open AI Gym, and stable_baselines3.
TABLE I: Performance of the proposed framework in health forecasting
@@ VI-A Predictive GraphRL Performance
Healthcare Forecasting: The proposed framework is evaluated to monitor health status by predicting future vital signs such as heart rate. Based on the sensor data and other clinical parameters such as ECG, Respiration, the time series prediction of the heart rate is conducted. The predicted values of heart for the next one hour are break-down into different time intervals (15 minutes, 30 minutes, 45 minutes, 60 minutes). Each of these time interval values acts as an observation for the GraphRL agent to monitor and communicate with the appropriate emergency team. The observation space of the vital sign, action space of different emergency teams and rewards for the agent actions in the predictive GraphRL environment are defined based on the modified early warning scores (MEWS) [38]. For the evaluation process, the WESAD dataset is adopted to conduct time series forecasting of heart rate. The proposed T-GCN in the predictive GraphRL environment performs better than the other baseline frameworks ELMA, GRU, and GRU-Based Multi-Agent as shown in Tab. I. It achieves the lowest MAE, MAPE and RMSE values in all the time intervals.
TABLE II: Performance of the proposed framework in traffic forecasting
@@ 15 Min 30 Min 45 Min 60 Min
Traffic Forecasting: The goal of the proposed framework is to predict traffic using the predictive GraphRL environment. The system takes in data with the following features: EventId, Type, Severity, TMC, Description, StartTime, EndTime, TimeZone, LocationLat, LocationLng, Distance, AirportCode, Number, Street, Side, City, County, State, and ZipCode. The observation space includes the current traffic state, which is represented by the traffic events and their severity in a particular region. The actions referred to possible traffic management strategies, such as altering traffic light timings or changing the speed limit. The rewards are defined based on the efficiency of the chosen strategy, such as reduced travel time or decreased congestion. For all the baseline models and the proposed framework, the MAE, MAPE, and RMSE values are reported for forecasting at 15, 30, 45, and 60-minute intervals. As shown in Tab. II, T-GCN outperforms the other models for all the forecasting intervals with the lowest MAE, MAPE, and RMSE values. The second-best performer is the GRU-Based Multi-Agent model, followed by GRU and ELMA.
TABLE III: Performance of the proposed framework in weather forecasting
@@
Weather Forecasting: In weather forecasting, the goal of the proposed framework is to use past weather data to predict future weather events and to optimise actions based on those predictions. In the predictive environment, the observation space is configured based on both the traffic and weather events datasets, including the event type, severity, start time, end time, location (latitude and longitude), and timezone. The actions represent the decisions the RL agent can take based on the observation space. For example, the agent could decide to issue a warning or alert for severe weather, adjust traffic signals or road signs, or change the speed limit on certain roads. The agent could receive a reward for correctly predicting severe weather and issuing a timely warning. Using the proposed GraphRL framework allows modelling the relationships between different weather events and their impact on traffic in a more efficient way than traditional machine learning methods. The GraphRL agent learns from these relationships to make better decisions and improve its predictions over time. Comparing the different models, T-GCN had the best performance across all metrics and different time intervals: 15, 30, 45, and 60 minutes, followed by GRU-Based Multi-Agent, GRU, and ELMA. The results show that the T-GCN model had the lowest MAE, MAPE, and RMSE values for all forecasting intervals, indicating its superior forecasting performance compared to the other models as shown in Tab. III.
@@ VI-B GraphRL Agent Performance
TABLE IV: Proposed GraphRL Performance
@@
The proposed RL agent was enabled with T-GCN and its performance is compared with other traditional RL agents as shown in Tab. IV. The table provides a comparison of different AI agents and their performance on three different datasets: WESAD, LAM Traffic Forecasting, and US Weather Forecasting. The performance of each agent is measured by a score, which is the total score achieved by the agent on the task over ten episodes. From the table, it can be seen that the proposed GraphRL agent is the most efficient agent on the WESAD dataset, as it scored the highest score. The DDPG and Q-Learning agents have the second-highest score on the WESAD dataset. On the LAM Traffic Forecasting dataset, the Q-Learning agent scored the lowest, and the proposed GraphRL agent scored the highest. On the US Weather Forecasting dataset, the A2C agent scored the lowest, while the GraphRL agent scored the highest. The GraphRL agent has outperformed other RL agents in all three predictive and monitoring applications.
[ xlabel=Episodes, ylabel=Score, legend pos=south east, legend style=font=, yticklabel style=/pgf/number format/fixed, ymin=-20000, ymax=65000, width=height=6cm, xtick=1,2,3,4,5,6,7,8,9,10, xticklabels=1,2,3,4,5,6,7,8,9,10]
[color=blue, mark=*] coordinates (1, 38925) (2, 43595) (3, 46875) (4, 57280) (5, 32245) (6, 485) (7, 48035) (8, 48790) (9, 42885) (10, 54470) ;
[color=red, mark=square*] coordinates (1, -8810) (2, -8100) (3, -11845) (4, 31960) (5, 46295) (6, 32560) (7, 53640) (8, 36195) (9, 36480) (10, 46425) ;
[color=green, mark=triangle*] coordinates (1, 23780) (2, -6765) (3, 33480) (4, 43635) (5, 24925) (6, 2670) (7, 8990) (8, 53145) (9, 58530) (10, 41405) ;
Health, Traffic, Weather

### v2

@@ IV-C GraphRL Agent
In this study, the Deep Q-Networks (DQN) algorithm is used. The Deep Q-Networks (DQN) algorithm, developed by Google’s DeepMind, was initially designed for playing Atari games. This algorithm enabled the AI to learn game strategies directly from visual input, without requiring pre-programmed rules or prior game-specific training. In this algorithm, the Q-Learning functions are approximated using the proposed T-GCN model, and the learning agent is rewarded based on the graph network prediction of the right action for the current state.
@@ IV-C1 Q-Function Approximation
T-GCN model used in this study to approximate the Q-Function for each action in the action space as shown in Fig. 2. The model is configured with parameters such as the relu activation function, mean square error as loss function, and Adam optimiser. The model gets trained with the state and its corresponding action. The learning agent performs an action $a_{t}\epsilon A$ for a transition from state $s_{t}$ to $s_{t}^{{}^{\prime}}$ and achieves a reward $R$ for the action. In this transition process, the maximum of the Q-function in Eq. 7 is calculated, and the discount of the calculated value uses a discount factor $\gamma$ to suppress future rewards and focus on immediate rewards. The discounted future reward is added to the current reward to get the target value. The difference between the current prediction from the neural networks and the calculated target value provides a loss function. The loss function is a deviation of the predicted value from the target value and it can be estimated from Eq. 8. The square of the loss function allows for the punishment of the agent for a large loss value.
@@ IV-C3 GraphRL Agent Algorithm
4:if $np.random.rand()<\epsilon$ then $\triangleright$ Exploration
@@
13:       $target\leftarrow R+\gamma\cdot\max(model.predict(s_{t+1}))$
@@
19:if $\epsilon\geq\epsilon_{min}$ then
20:    $\epsilon\leftarrow\epsilon\cdot\epsilon_{decay}$
@@
Algorithm 2 introduces the GraphRL Agent, presents the functionality of the GraphRL Agent within a complex action-state environment, utilizing T-GCN for Q-function approximation. This integration enables effective handling of spatial-temporal data complexities, enhancing decision-making. The algorithm’s design is founded on a strategic balance between exploration and exploitation, achieved via an epsilon-greedy strategy, crucial for adaptive learning and continual improvement in decision-making. It starts with initializing key parameters (Line 1), defining a T-GCN model for handling complex data structures (Line 2), and storing memories for experience replay (Line 3). The agent’s learning process involves iterative learning from minibatches of experiences, computing target Q-values, and adjusting its policy (Lines 8-21), with a dynamic adjustment of the exploration rate (Lines 16-18). The predict() function, pivotal in the exploitation phase, utilizes the T-GCN model’s predictions to guide actions, showcasing the algorithm’s advanced approach in navigating dynamic environments through a blend of exploration and strategic exploitation.
@@ IV-C4 Implementation Algorithm
2:$\mathcal{C}=\{1,2,\ldots,C\}$: set of subjects
3:$\mathcal{V}=\{1,2,\ldots,V\}$: set of vital signs
4:=$\mathcal{M}=\{1,2,\ldots,M\}$: number of episodes
@@
Algorithm 3 serves as the comprehensive implementation of the GraphRL framework, intricately combining the Predictive GraphRL Environment (Algorithm 1) with the GraphRL Agent (Algorithm 2). This pivotal algorithm orchestrates the real-time interactions between the agent and the environment, thus forming the operational core of the framework. It outlines the simulation scope, input parameters (subjects $\mathcal{C}$, vital signs $\mathcal{V}$, and episodes $\mathcal{M}$), and the output in the form of cumulative rewards. The initialization phase prepares the environment env and the agent for interaction. The episodic loop, encompassing the agent’s action-response cycle, is vital for continuous learning and adaptation. Crucial to this process is the memorize() function, which stores experiences for later recall during experience replay, allowing the agent to learn from past actions and refine its decision-making strategy. This algorithm thus encapsulates the dynamic and iterative nature of the GraphRL framework, highlighting the importance of memory and experience in the realm of advanced reinforcement learning, and showcasing its functionality in complex, evolving environments.
@@ IV-D Bayesian Optimisation
Bayesian optimisation is a global optimisation method that uses a probabilistic model to guide the search for the optimal solution. The model is updated as new data points are sampled and evaluated, allowing the algorithm to improve its predictions over time by fine-tuning the $L_{reg}$, $\lambda$ and minimising the loss function defined in Eq. 6. The basic idea behind Bayesian optimisation is to model the objective function, f(x), as a Gaussian process (GP). The GP model is used to predict the objective function value at any point x, given the observations of the function at other points. The prediction is given by the posterior distribution of the GP, which is a Gaussian distribution with mean and variance given by Eq. 9.
@@
Where X is the matrix of previously sampled points, y is the vector of corresponding function values, K is the Gram matrix of the covariance function evaluated at X, and $\sigma^{2}$ is the noise level in the function evaluations. The next point to sample is chosen based on an acquisition function, which balances the trade-off between exploration and exploitation. Common acquisition functions include the probability of improvement and the expected improvement. Given a set of observed points (X,y) and a Gaussian process prior, Bayesian optimisation seeks the point x* that minimises the loss function value, given by Eq. 10. The optimisation process continues iteratively, sampling new points and updating the GP model until a stopping criterion is met.
@@ V-A Datasets
The GraphRL framework’s testing with datasets from healthcare, traffic, and weather domains was a deliberate strategy to assess its versatility and robustness in handling diverse time-series data. The choice of these varied domains was intended to demonstrate the framework’s adaptability and efficacy in different contexts. Each domain poses unique challenges: healthcare data’s complexity and sensitivity, traffic data’s dynamic patterns requiring real-time analysis, and weather data’s intricate interplay of environmental factors. Successfully navigating these distinct datasets underscores the framework’s capability for widespread real-world application. Additionally, using datasets from different fields facilitates a thorough evaluation of the framework, ensuring its versatility and effectiveness across various problem types and data structures. This comprehensive approach is vital for a tool designed for extensive applications in data analysis and prediction. Three datasets utilized for evaluating the GraphRL framework as shown in Fig. I, each from a different domain: healthcare, traffic, and weather. In healthcare, the WESAD dataset, containing electrocardiogram (ECG) and photoplethysmogram (PPG) data from 17 participants, offers a rich source of biometric time-series data for pattern recognition analysis. The Los Angeles (LA) Traffic dataset, sourced from the Los Angeles Department of Transportation (LADOT), provides real-time urban traffic data like traffic counts and speeds, while the Large-Scale Traffic and Weather Events (LSTW) dataset, with data across the United States, uniquely combines traffic conditions and weather events, posing a multifaceted challenge for the framework.
TABLE I: Datasets
Dataset Domain Key Features Statistics Suitability
WESAD
[26]
Healthcare
Physiological data
(ECG, PPG),
motion data
17 participants
Rich biometric time-series
data, ideal for testing pattern
recognition in health-related
data.
LA
Traffic
[27]
Traffic
Traffic counts,
speeds,
travel times
Data from
LADOT
Real-time urban traffic data,
useful for analyzing and
predicting dynamic traffic flows.
LSTW
[28]
Weather
Traffic conditions,
weather events
Data across
the United
States
Combines traffic and weather
data, challenging the framework
to handle complex, multifactorial
scenarios.
@@ V-B Baseline Models
In our study, we selected three baseline models for comparison, each epitomizing state-of-the-art approaches in multi-agent forecasting, graph neural networks, and traffic prediction. These models were chosen based on their innovative methodologies and proven effectiveness in areas closely aligned with our research objectives.
  • ELMA Method [29]: Developed by Li et al. [29], the ELMA method utilizes graph neural networks for forecasting multi-agent activities, particularly adept at handling spatiotemporal data. Its novelty lies in the use of energy-based learning, making it an excellent benchmark against our framework, which similarly leverages graph-based techniques in complex environments.
  • Self-Supervised Technique [30]: This technique is pioneering in self-supervised learning for predicting multi-agent driving behavior. Its relevance to our study comes from its focus on behavior prediction in diverse scenarios, using self-supervised domain knowledge—an advanced trend in multi-agent learning.
  • Internet Traffic Prediction Study [31]: It involves internet traffic prediction using distributed multi-agent learning, employing LSTM and GRU models. GRU’s superior performance in their study provides a valuable point of comparison for our research, which focuses on sophisticated learning techniques in traffic prediction.
Each of these models represents a significant stride in their respective fields. Their selection for comparison in our study is justified by their alignment with our research goals and their benchmark status in handling complex, dynamic datasets. By comparing our GraphRL framework against these models, we aim to demonstrate our approach’s novelty and effectiveness in diverse real-world applications.
@@ V-C Evaluation Metrics
Mean Absolute Error (MAE) is a widely-utilized regression metric that gauges the average magnitude of errors between predicted and actual values in a dataset. It is calculated by averaging the absolute differences between these values, yielding a singular metric. Root Mean Squared Error (RMSE) is another prominent regression metric, assessing the average magnitude of differences between predicted and actual values. RMSE is computed as the square root of the mean of these squared differences. Mean Absolute Percentage Error (MAPE) represents yet another regression metric, quantifying the average absolute percentage error between predicted and actual values. It is derived by averaging the absolute differences between these values, expressed as a percentage of the actual values. Conversely, Cumulative Rewards is a performance metric specific to reinforcement learning. It measures the total rewards an agent accumulates over a specified timeframe or across a set number of actions, calculated by summing all rewards received during this period.
In the context of the experiments conducted for this study, Python version 3.7.6 served as the programming environment, with the deployment of several packages including TensorFlow, Keras, OpenAI Gym, and stable_baselines3.
TABLE II: Performance of the proposed framework in health forecasting
@@ VI-A Predictive GraphRL Performance
Healthcare Forecasting: The proposed framework is evaluated to monitor health status by predicting future vital signs such as heart rate. Based on the sensor data and other clinical parameters such as ECG, Respiration, the time series prediction of the heart rate is conducted. The predicted values of heart for the next one hour are break-down into different time intervals (15 minutes, 30 minutes, 45 minutes, 60 minutes). Each of these time interval values acts as an observation for the GraphRL agent to monitor and communicate with the appropriate emergency team. The observation space of the vital sign, action space of different emergency teams and rewards for the agent actions in the predictive GraphRL environment are defined based on the modified early warning scores (MEWS) [32]. For the evaluation process, the WESAD dataset is adopted to conduct time series forecasting of heart rate. The proposed T-GCN in the predictive GraphRL environment performs better than the other baseline frameworks ELMA, GRU, and GRU-Based Multi-Agent as shown in Tab. II. It achieves the lowest MAE, MAPE and RMSE values in all the time intervals.
TABLE III: Performance of the proposed framework in traffic forecasting
@@ 15 Min 30 Min 45 Min 60 Min
Traffic Forecasting: The goal of the proposed framework is to predict traffic using the predictive GraphRL environment. The system takes in data with the following features: EventId, Type, Severity, TMC, Description, StartTime, EndTime, TimeZone, LocationLat, LocationLng, Distance, AirportCode, Number, Street, Side, City, County, State, and ZipCode. The observation space includes the current traffic state, which is represented by the traffic events and their severity in a particular region. The actions referred to possible traffic management strategies, such as altering traffic light timings or changing the speed limit. The rewards are defined based on the efficiency of the chosen strategy, such as reduced travel time or decreased congestion. For all the baseline models and the proposed framework, the MAE, MAPE, and RMSE values are reported for forecasting at 15, 30, 45, and 60-minute intervals. As shown in Tab. III, T-GCN outperforms the other models for all the forecasting intervals with the lowest MAE, MAPE, and RMSE values. The second-best performer is the GRU-Based Multi-Agent model, followed by GRU and ELMA.
TABLE IV: Performance of the proposed framework in weather forecasting
@@
Weather Forecasting: In weather forecasting, the goal of the proposed framework is to use past weather data to predict future weather events and to optimise actions based on those predictions. In the predictive environment, the observation space is configured based on both the traffic and weather events datasets, including the event type, severity, start time, end time, location (latitude and longitude), and timezone. The actions represent the decisions the RL agent can take based on the observation space. For example, the agent could decide to issue a warning or alert for severe weather, adjust traffic signals or road signs, or change the speed limit on certain roads. The agent could receive a reward for correctly predicting severe weather and issuing a timely warning. Using the proposed GraphRL framework allows modelling the relationships between different weather events and their impact on traffic in a more efficient way than traditional machine learning methods. The GraphRL agent learns from these relationships to make better decisions and improve its predictions over time. Comparing the different models, T-GCN had the best performance across all metrics and different time intervals: 15, 30, 45, and 60 minutes, followed by GRU-Based Multi-Agent, GRU, and ELMA. The results show that the T-GCN model had the lowest MAE, MAPE, and RMSE values for all forecasting intervals, indicating its superior forecasting performance compared to the other models as shown in Tab. IV.
@@ VI-B GraphRL Agent Performance
TABLE V: Proposed GraphRL Performance
@@
The proposed RL agent was enabled with T-GCN and its performance is compared with other traditional RL agents as shown in Tab. V. The table provides a comparison of different AI agents and their performance on three different datasets: WESAD, LAM Traffic Forecasting, and US Weather Forecasting. The performance of each agent is measured by a score, which is the total score achieved by the agent on the task over ten episodes. From the table, it can be seen that the proposed GraphRL agent is the most efficient agent on the WESAD dataset, as it scored the highest score. The DDPG and Q-Learning agents have the second-highest score on the WESAD dataset. On the LAM Traffic Forecasting dataset, the Q-Learning agent scored the lowest, and the proposed GraphRL agent scored the highest. On the US Weather Forecasting dataset, the A2C agent scored the lowest, while the GraphRL agent scored the highest. The GraphRL agent has outperformed other RL agents in all three predictive and monitoring applications.
[…]